A particle is executing Simple Harmonic Motion of amplitude 6 m and period 3× 5

seconds. Find the maximum velocity of the particle.


1
Expert's answer
2021-06-01T11:24:35-0400

"A=6\ m, T=3.5\ s"



"\omega=\dfrac{2\pi}{T}=\dfrac{2\pi}{3.5}rad/s""v_{max}=A\omega=6(\dfrac{2\pi}{3.5}) \ m/s=\dfrac{24\pi}{7}\ m/s"

The maximum velocity of the particle is "\dfrac{24\pi}{7}" m/s.